Potassium ferrocyanide is 50% ionized in aqueous
solution, its van't Hoff factor will be equal to ——.

Grade:12

2 Answers

vikas askiitian expert
13 years ago

K4[Fe(CN)6]•3H2O => 4K+ + [Fe(CN)6]4-

 

1                               0             0                          (initial)

1-x                            4x          x                         (atfter dissociation)

 

vant hoff factor (i) = moles after dissotion/total initial moles

                          = 1+4x

50% ionise means x = 0.5 so

 i = 1 + 4*0.5 = 3

its vant hoff factor is 3

Yash Chourasiya
3 years ago
Hello Student

K4[Fe(CN)6].3H2O→ 4K++ [Fe(CN)6]4-
1 0 0 (initial)

(1-x) 4x x (atfter dissociation)

vant hoff factor (i) = moles after dissotion/total initial moles
= 1+4x
50% ionise means x = 0.5
so, i = 1 + 4*0.5 = 3

Vant hoff factor is 3

I hope this answer will help you.